# Unleashing Large Multimodal Model’s Ability for Painting 3D Objects with Realistic Materials

### v2

@@ 3.6 SVBRDF Maps Generation
Figure 5: Qualitative results of Make-it-Real refining 3D asserts without PBR map. Objects are selected from Objaverse [14] with diffuse only.
Figure 6: Qualitative results of Make-it-Real refining 3D object generated by edge-cutting TripoSR [50] conditioned on single image. The first line per two are initial objects generated by TripoSR, while the second line are objects refined by Make-it-Real.
@@
Figure 7: Qualitative results of Make-it-Real refining 3D object generated by edge-cutting text-to-3D models.
Consequently, we successfully develop a pipeline for generating the material maps from 3D objects with constrained diffuse maps. This process finally project the generated material textures back onto the 3D meshes, thereby achieving a natural and authentic appearance, along with the physical properties of the materials available for photo-realistic rendering in different environments.
@@ 4 Experiments
To verify the effectiveness of the proposed method, we conduct refinement experiments mainly on two types of objects. The first type is artificial 3D models, with the primary model from Objaverse [14]; The second type is objects generated by state-of-the-art 3D generation methods. We validate the powerful capabilities of our method through extensive qualitative and quantitative experiments on both types of 3D models.
4.1 Texture refinement for existing 3D assets
We observe that the realism of objects is significantly impacted by the existing publicly available 3D assets, which have only diffuse maps and lack material properties. This deficiency extends to artificial 3D models derived from Computer-Aided Design (CAD), which necessitate automated texture assignment. To address this issue, we conduct texture synthesis on 3D models from Objaverse [14], which are available only with diffuse maps. Our methodology, Make-it-Real, is employed to generate additional material texture maps. The qualitative outcomes are illustrated in  Fig. 5. Make-it-Real demonstrates the capability to accurately segment objects, assign multiple appropriate materials, and synthesize textures with high fidelity and photorealism.
Assets enhanced by Make-it-Real exhibit the capability for relighting operations, with some materials displaying significant highlights, such as a bathtub made of marble, a globe constructed from gold, and a vase crafted from ceramic. These different materials interact with light to varying degrees, resulting in distinct reflections of the same environment light and presenting varied textures. Moreover, the material characteristics change across regions; for example, the land and handle parts of the globe are imbued with the properties of gold, while other parts are identified as plastic. In addition, the material can alter in terms of roughness and appearance, such as the fine wrinkles that appear on the red box in the second column, and the blue box’s base showing more pronounced color contrast, enhancing realism and reflecting the object’s age of use.
4.2 Texture refinement for generated 3D object
Domain Method Base object +Make-it-Real
@@ 4.2 Texture refinement for generated 3D object
Table 1: GPT evaluation. GPT’s preference comparison on Make-it-Real refined objects generated by state-of-the-art 3D generation methods with their original objects.
Figure 8: Ablation study for texture refinement module.
Figure 9: Ablation study for effect of different queried maps.
We test the effectiveness of the proposed pipeline for texture refinement on generated 3D models. For image-to-3D, we use the state-of-the-art model TripoSR [50]. For text-to-3D, we use objects generated by typical models like Instant3D, MVDream and Fantasia3D [30, 46, 9]. Make-it-Real is used to refine these generated objects. Figure 6 displays the qualitative results of Tripo-SR, with the referenced image on the left. We note that the Tripo-SR model is capable of generating accurate shapes and consistent diffuse maps, yet it struggles to reconstruct material effects suitable for relighting in the images. By leveraging our model for enhancement, we observe that Make-it-Real successfully generates appropriate material maps for these models. In the first image, the badge exhibits a metallic sheen, and the clothing possesses a more fabric-like texture; in the second image, both the robot’s eyes and body parts exhibit brightness. The effects become more pronounced under different HDR conditions, demonstrated in the appendix. Figure 7 illustrates the outcomes of three different text-to-3D models. Similarly, our Make-it-Real successfully paint these models with materials.
We employ the same idea of GPTEval3D [56] for fair, human-aligned assessment, utilizing the powerful GPT4V model to compare models before and after refinement. Results are shown in  Tab. 1. Compared to the original generated 3D models, our refined models are capable of enhancing the realism of 3D objects across all edge-cutting methods for 3D generation. Notably, as the 3D generation methods evolve and prior knowledge increases (from text prompts to single-view images), the realism of our refined results also improves. This validates that our approach can better endow materials with enhanced realism on the foundation of more advanced generation techniques.
Figures 6 and 7 display the qualitative experimental results. Figure 6 showcases objects refined based on TripoSR [50]. Our Make-it-Real performs reasonable segmentation and material matching on object parts, refining them into more realistic models. Notably, as image-to-3D setting can utilize a high-resolution single-view image as prior, we first generate a detailed description of these images using MLLM. This can compensate for the lack of material information in images rendered from the initial 3D object (as seen in the first example of the police metal badge), thereby generating better results. Figure 7 presents the refinement results on different text-to-3D methods. Compared to image-to-3D, the original quality of these objects is relatively lower, yet our method still robustly matches the correct materials, thereby enhancing the quality of the objects.
@@ 4.3 Ablation Study
Missing Part Refinement. Upon completing the query retrieval steps, we precisely assign material types to the segmented regions within the view, thereby initially determining the material categories for each block on the texture map. However, due to limitations such as the number of views, there may exist regions on the texture map where materials are not assigned. To address this issue, we design a heuristic texture refinement post-processing operation. Specifically, we calculate the mean color values of the diffuse maps for regions where materials have been assigned and use this as a basis for clustering the remaining unassigned areas. This operation extends the material mapping of the texture map to full resolution, ensuring overall consistency in the texture of the 3D object while avoiding the absence of textures. This method not only improves the coverage of textures but also enhances the visual quality of the final 3D model. Results are shown in Fig. 8.
Effects of different material maps. To examine the individual contributions of various material maps to the rendering process, we incrementally introduced each material map—specifically, roughness, metalness, and displacement maps—into our model, shown in Fig. 9. The observed effects of these incremental additions were systematically analyzed. For instance, the inclusion of the roughness map was found to significantly enhance the depiction of surface textures, imparting a noticeable wrinkled appearance, while the introduction of the metalness map resulted in enhanced reflective properties, contributing to the realism of metallic surfaces. Furthermore, the combination of both roughness and displacement maps was demonstrated to produce a pronounced three-dimensional effect, as illustrated in the second row of our results. These outcomes emphatically validate the utility of querying each specific material map within our framework. Our pipeline is designed to ensure the retrieval of a comprehensive set of material attributes, thereby facilitating their effective application in downstream 3D rendering pipelines.
5 Conclusion and Discussion
@@
Limitations and future work. While promising results have been achieved by MLLMs for texture assignment through Make-it-Real. Our work still faces several unresolved challenges. First, although our method can achieve appropriate texture assignment for diffuse-only model, it does not support reverse transform from shaded texture map to diffuse map for generated 3D objects. This causes problem of assigning different materials to the dark shadow and highlight area when generated object with mesh already shaded in different lighting conditions. Second, we find base model quality is essential for MLLM to assign correct materials. When base 3D object is of low quality, it is difficult for MLLMs to identify object properties when ground truth text prompt describing the object is not available. We consider methods mitigating these challenges as well as adding user friendly control into our fully-automatic pipeline as future work.
Appendix 0.A Supplementary Material Overview
@@
Appendix 0.B Details of Make-it-Real
In this section, we detail the pipeline that briefly outlined in the main paper. We commence by elaborating on the generation of SVBRDF Maps, incorporating illustrative figures to detail the process involving operations in computer graphics. Sec. 0.B.1 is dedicated to explaining the acquisition of region-level texture map partition. Following that, in Sec. 0.B.2 we discuss the method of pixel-level diffuse-referenced estimation. Then, in Sec. 0.B.3 we declare some details of rendering procedure. Sec. 0.B.4 details the prompt design for material captioning and matching. Sec. 0.B.5 reports the details of the GPT-4V based evaluation.
0.B.1 Texture Partition Module Design
As illustrated in Fig. 10, our process initiates with the rendering of a 3D object incorporating the original diffuse (i.e. query diffuse) from multiple perspectives. Following this rendering phase, we employ GPT-4V alongside a segmentor to derive segmented masks for the materials associated with each viewpoint. The subsequent step involves the extraction of regions masked in these images and their back-projection onto the mesh of the object. By examining the object from all acquired viewpoints and applying UV unwrapping techniques, we achieve preliminary segmentation of all materials. Subsequently, each material’s segment is refined using a diffuse-based mask refinement operation. Ultimately, by combining the segments of all materials, we obtain a region-level texture partition map, which serves to guide our subsequent, more detailed operations.
@@ 0.B.1 Texture Partition Module Design
Figure 11: Pixel-level diffuse-referenced estimation module. We generate spatially variant BRDF maps by referencing diffuse maps, employing KD-Tree algorithm for efficient nearest neighbor searches, and normalizing colors via histogram equalization.
0.B.2 Diffused-referenced Module Design
@@
As illustrated in Fig. 11, we developed a pixel-level diffuse-referenced estimation module, building upon the foundations laid out in Sec. 0.B.1. This module is inspired by a technique frequently employed by 3D artists, who often utilize diffuse maps as a reference to generate images of other material properties. Accordingly, we designate the known diffuse map as the query diffuse, and the diffuse corresponding to the region of interest in the material as the key diffuse.
@@
The process to precisely obtain the final SVBRDF maps is divided into four steps: 1) Initially, to address potential gaps in color intensity between the two diffuse maps, histogram equalization is employed to achieve a more uniform color distribution across the image. 2) Subsequently, for each pixel on the query diffuse—termed a reference pixel—we seek the most similar neighboring color index on the key diffuse. Given the high dimensionality of both maps (typically 1024x1024 pixels), a brute-force approach to this search would be computationally prohibitive. To this end, we accelerate the pixel query process using a KD-Tree algorithm, which organizes the RGB values of the key map’s diffuse into a KD-Tree for efficient nearest neighbor searches, reducing the computation time to under ten seconds. 3) The third step involves using the obtained indices to obtain corresponding values from the rest of the material maps. 4) Finally, by aggregating the query results for all material segments, we are able to generate the comprehensive spatially variant SVBRDF maps.
0.B.3 Rendering Procedure Details
@@
In the context of computer graphics, “diffuse” typically denotes the primary color of a material, a concept analogous to “base color” in Physically-Based Rendering (PBR) paradigms, both representing the inherent color of the material under uniformly scattered illumination.
We observe that in the provided material collection [52] that a portion of the materials exhibit either completely black or low-quality diffuse maps. Therefore, for those materials with missing or subpar diffuse maps, we pragmatically substitute the base color map as the key diffuse map. In the majority of instances, our rendering adheres to the Cook-Torrance model [13], predominantly employing the original diffuse map, alongside our generated spatially variant normal map, metalness map, and roughness map to render objects. Depending on the workflow requirements, the inclusion of a height map, displacement map, specular map, and combinations of additional maps is optional. Regarding 2D-3D alignment techniques (such as back-projection and UV unwrapping), we rely on the Kaolin [25] package and back-projection with masks [45] techniques, with model parameters maintained consistent with  [45]. This approach allows for the preservation of material visual quality in scenarios where the diffuse map is missing or of inferior quality, by leveraging the base color map as an effective substitute, thereby ensuring consistency and realism within the PBR workflow. Furthermore, the flexible application of various map types and 2D-3D alignment techniques enhances the detail and realism of rendered objects, meeting the demands of diverse rendering scenarios.
@@
0.B.4 Prompt Details for Multi-modal Large Language Models
@@
Material captioning. Due to limit context length for image tokens, it is currently impossible for MLLMs to directly memories thousands of images of material balls. To address this, we use GPT-4V to generate detailed captions for each material ball. As demonstrated in Fig. 12, we input concatenated images of material ball from same subcategory with prompt specifically tailored to highlight texture properties. This strategy guides GPT-4V to generate detailed caption for each material ball, distinguishing the subtle differences between them. These detailed captions are then registered into material library.
Material matching. The MLLM-based material query process is exemplified through a simplified case, as illustrated in Fig. 13. Initially, GPT-4V is queried to identify the basic material for matching. Following this preliminary matching, GPT-4V is queried for more specific materials using the names of subcategories as filters, which narrows down the selection to a few candidates within the same subcategory. For final material selection, GPT-4V is prompted with detailed captions from the library, directing it to allocate the most suitable material. This is accomplished in conjunction with a meticulously engineered segmentation process in UV space, ultimately facilitating MLLM-based material matching. It is noteworthy that prompting GPT-4V to navigate through a three-level tree structure has been proven to enhance both the efficiency and accuracy of the matching process, as opposed to directly selecting from thousands of materials directly.
@@
0.B.5 Evaluation Details of Quantitative Results
Data statistics. For TripoSR [50], we generate objects conditioned on 10 images from its official code-base and 100 images rendered from randomly sampled objaverse [15] objects. For  [30, 46], we use text prompts proposed in GPTeval3D [56] to generate 130 objects. For Fantasia3D [9], we use 15 text prompts from it’s original code-base and 15 text prompts designed by our human experts. Fantasia3D generate objects less than other methods due to the necessity for meticulous hyper-parameter design to generate meaningful objects.
Evaluation detail. For the purpose of evaluating objects pre- and post-refinement, we render four view images of each object and concatenate them vertically to facilitate a comprehensive assessment. We craft a specific prompt to guide GPT-4V in conducting an impartial comparison of the texture quality between two objects. As demonstrated in Fig. 14. GPT-4V’s advanced capabilities enable it to differentiate between the two objects, providing comparison results that closely align with assessments made by human experts. In the scenario evaluated, GPT-4V successfully identifies all enhancements (metal badges, human skin, and fabric textures).
Appendix 0.C Additional Related Works
Figure 15: Comparison between previous method and Make-it-Real. We demonstrate the distinctions between Material Palette[35] and our method in terms of material identification and extraction. Our overall pipeline presents a more challenging task, where the input is a rendered image with only diffuse information, and the output consists of textures for the entire object.
@@ Appendix 0.C Additional Related Works
Our problem setting poses a more challenging task, as illustrated in Fig. 15, which involves identifying and recovering the original material properties from images of objects that only have a diffuse map, in addition to generating different material maps for the entire object. Humans have the capacity to intuitively infer the underlying material properties of 3D objects from single images containing only shape and basic colors. This capability is attributed to our robust material recognition abilities and comprehensive prior knowledge of object categories, colors, and material attributes. Building on this concept, our approach harnesses the potent image recognition capabilities and prior knowledge inherent in large-scale multimodal language models to efficiently execute region-level material extraction and identification, which is further enhanced by subsequent 2D-3D alignment and diffuse-reference techniques to generate and apply material texture maps for physically realistic rendering of 3D objects.
Appendix 0.D Additional Experiments
0.D.1 User Preference Study
We randomly sample 100 objects generated from each methods (30 for Fantasia3D [9]) from Sec. 0.B.5 for human volunteers. Their task is to express their preference for objects either before or after undergoing refinement by Make-it-Real. To facilitate this evaluation, we concatenate images rendered from four views of each object, both pre- and post-refinement, in a random sequence. Volunteers are instructed to select the version they preferred, based on texture quality and photo-realism. The statistical outcomes of this study are depicted in Fig. 16. It is observed that volunteers consistently favor the objects post-refinement across all evaluated 3D content creation methods. This preference aligns with the evaluations performed by GPT-4V, indicating a general consensus on the enhancement in quality achieved through Make-it-Real refinement process.
Figure 16: User preference study of pre- and post-refinement of objects generated by edge-cutting 3D generation methods.
0.D.2 Effects of Different Texture Maps

### v3

@@ 3.2.3 SVBRDF Maps Generation
Figure 5: Qualitative results of Make-it-Real refining 3D asserts without PBR maps. Objects are selected from Objaverse [14] with albedo maps only.
Figure 6: Visualization of generated texture maps. We visualize some SVBRDF maps, where the material maps are well aligned with the albedo maps.
@@
Figure 7: Qualitative comparisons between Make-it-Real refining results and 3D objects generated by edge-cutting 3D content creation models. The upper row depicts image-to-3D models (InstantMesh and TripoSR), and the lower row shows results of text-to-3D models.
@@ 4 Experiments
4.1 Experiment settings
To verify the effectiveness of Make-it-Real, we conduct refinement experiments mainly on two types of objects. The first type is artificial 3D assets, with the primary model from Objaverse [14]; The second type is objects generated by state-of-the-art 3D generation methods. For existing 3D assets, we pick 200 objects with diverse textures from Objaverse by human experts. For 3D generative models (InstantMesh [63], TripoSR [53], MVDream [49], Instant3D [32] and Fantasia3D [9]), we also generate 200 objects for each methods use prompts designed in GPTEval3D [62]. We use Make-it-Real to refine the objects and compare the texture quality before and after refinement. We perform both GPT-4V [42] based evaluation and user study on the above objects (Detailed guidence and prompts are available in  Section B.5).
4.2 Experiment results
Texture refinement for existing 3D assets. As shown in Figure 5, assets processed through Make-it-Real demonstrate the capability to accurately segment objects, assign various suitable materials, and synthesize high-fidelity, photorealistic textures. Some materials exhibit notable highlights, such as a marble bathtub and a globe made of gold. The interaction of these different materials with light varies significantly, leading to diverse reflective effects under the same environmental lighting, thus presenting a range of textures. Additionally, material properties vary across different regions; for instance, the globe’s landmass and handle exhibit gold characteristics, while other parts are identified as plastic. Furthermore, the texture and appearance of materials also vary, such as the subtle wrinkles on the red box in the second column and the more pronounced color contrast at the base of the blue box, which enhances realism and reflects signs of use. Due to the albedo-referenced algorithm design in Section 3.2.3, regions with similar colors have similar BRDF values, avoiding abrupt changes in material properties, such as the continuous gold surface of the globe and the silver body of the kettle. Regions with significant color differences display consistent differential distributions with the key map, such as the embossed textures on the lower left corner of the water bottle in Figure 5 and the subtle particle variations on the surface of the red oil drum. Additionally, visualization of texture maps demonstrates reasonably consistent textures with albedo, shown in Figure 6. Quantitive results are shown in Table 1, and more qualitative results can be found in Appendix E.
Texture refinement for generated 3D objects. Figure 7 displays the qualitative results. By leveraging our model for enhancement, we observe that Make-it-Real successfully generates appropriate material maps for both image-to-3D and text-to-3D models.
Similarly, our Make-it-Real successfully paint these models with materials. As reported in Table 1, human experts consistently favor the objects post-refinement across all evaluated 3D content creation methods. This preference aligns with the evaluations performed by GPT-4V, indicating a general consensus on the enhancement in quality achieved through Make-it-Real refinement process.
Table 1: GPT evaluation and user preference. GPT’s and user’s preference comparison on Make-it-Real refined objects sourced from existing 3D assets and state-of-the-art 3D generation methods.
Domain Method / Source GPT Evaluation User Preference
Base object +Make-it-Real Base object +Make-it-Real
@@ 4.2 Experiment results
[…]
Figure 8: Ablation study of material segmentation refinement. Compared to direct usage of SemanticSAM [31], Our post-process tailored for material segmentation on 3D object can produce more consistent results.
Figure 9: Ablation study of missing part refinement. Our method on the bottom row produces consistent texture maps and avoids the missing parts of material texture.
@@ 4.3 Ablation Study
Effects of Mask Refinement Module. In Section 3.2.1, we performed additional material post-processing on the segmentation outcomes, as depicted in Figure 9. The refined results indicate that our module achieves precise material segmentation for most standard objects, thereby enabling more accurate queries by GPT-4V. In Section 3.2.3, we addressed the completion of missing regions in the texture map within the UV space at the regional level, as illustrated in Figure 9. This method not only increases texture coverage but also enhances the visual quality and consistency of the refined 3D model. More ablation studies can be found in Section D.1.
5 Conclusion
@@
Appendix A Supplementary Material Overview
@@
Appendix B Details of Make-it-Real
In this section, we detail the pipeline that briefly outlined in the main paper. We commence by elaborating on the generation of SVBRDF Maps, incorporating illustrative figures to detail the process involving operations in computer graphics. Section B.1 is dedicated to explaining the acquisition of region-level texture map partition. Following that, in Section B.2 we discuss the method of pixel-level albedo-referenced estimation. Then, in Section B.3 we declare some details of rendering procedure. Section B.4 details the prompt design for material captioning and matching. Section B.5 reports the details of the GPT-4V based evaluation. Potential broader impacts are discussed in Appendix F.
B.1 Texture Partition Module Design
As illustrated in Figure 10, our process initiates with the rendering of a 3D object incorporating the original albedo (i.e. query albedo) from multiple perspectives. Following this rendering phase, we employ GPT-4V alongside a segmentor to derive segmented masks for the materials associated with each viewpoint. The subsequent step involves the extraction of regions masked in these images and their back-projection onto the mesh of the object. By examining the object from all acquired viewpoints and applying UV unwrapping techniques, we achieve preliminary segmentation of all materials. Subsequently, each material’s segment is refined using a albedo-based mask refinement operation. Ultimately, by combining the segments of all materials, we obtain a region-level texture partition map, which serves to guide our subsequent, more detailed operations.
@@ B.1 Texture Partition Module Design
Figure 11: Pixel-level albedo-referenced estimation module. We generate spatially variant BRDF maps by referencing albedo maps, employing KD-Tree algorithm for efficient nearest neighbor searches, and normalizing colors via histogram equalization.
B.2 Albedo-referenced Module Design
@@
As illustrated in Figure 11, we developed a pixel-level albedo-referenced estimation module, building upon the foundations laid out in Section B.1. This module is inspired by a technique frequently employed by 3D artists, who often utilize albedo maps as a reference to generate images of other material properties. Accordingly, we designate the known albedo map as the query albedo, and the albedo corresponding to the region of interest in the material as the key albedo.
@@
The process to precisely obtain the final SVBRDF maps is divided into four steps: 1) Initially, to address potential gaps in color intensity between the two albedo maps, histogram equalization is employed to achieve a more uniform color distribution across the image. 2) Subsequently, for each pixel on the query albedo—termed a reference pixel—we seek the most similar neighboring color index on the key albedo. Given the high dimensionality of both maps (typically 1024x1024 pixels), a brute-force approach to this search would be computationally prohibitive. To this end, we accelerate the pixel query process using a KD-Tree algorithm, which organizes the RGB values of the key map’s albedo into a KD-Tree for efficient nearest neighbor searches, reducing the computation time to under ten seconds. 3) The third step involves using the obtained indices to obtain corresponding values from the rest of the material maps. 4) Finally, by aggregating the query results for all material segments, we are able to generate the comprehensive spatially variant SVBRDF maps.
B.3 Rendering Procedure Details
@@
In the context of computer graphics, “albedo” typically denotes the primary color of a material, a concept analogous to “base color” and “diffuse” in Physically-Based Rendering (PBR) paradigms, both representing the inherent color of the material under uniformly scattered illumination.
In accordance with workflow requirements, the inclusion of height maps, displacement maps, specular maps, and combinations of additional maps is optional. For objects demanding high surface details, such as the oil barrel in Figure 5 and the stone horse in Figure 16, displacement or height maps are incorporated and rendered using the mature 3D computer graphics software Blender engine. Regarding 2D-3D alignment techniques, including rasterization rendering, back-projection, and UV unwrapping detailed in Section 3.2.1, we adhere to the methodologies outlined in [47]. Specifically, we utilize the Kaolin package (Kaolin [26]) and back-projection with masks [47], with the model learning rate set to 0.02 in Equation 3 and the difference coefficient $\sigma$ set to 0.1 in Equation 4.
This approach facilitates the systematic generation of various material texture maps in the presence of only an albedo map, thereby ensuring consistency and realism within the Physically Based Rendering (PBR) workflow. Moreover, the flexible application of different map types and 2D-3D alignment techniques significantly enhances the detail and realism of rendered objects, effectively meeting the demands of diverse rendering scenarios.
@@
B.4 Prompt Details for Multi-modal Large Language Models
@@
Material captioning. Due to limit context length for image tokens, it is currently impossible for MLLMs to directly memories thousands of images of material balls. To address this, we use GPT-4V to generate detailed captions for each material ball. As demonstrated in Figure 12, we input concatenated images of material ball from same subcategory with prompt specifically tailored to highlight texture properties. This strategy guides GPT-4V to generate detailed caption for each material ball, distinguishing the subtle differences between them. These detailed captions are then registered into material library.
Material matching. The MLLM-based material query process is exemplified through a simplified case, as illustrated in Figure 13. Initially, GPT-4V is queried to identify the basic material for matching. Following this preliminary matching, GPT-4V is queried for more specific materials using the names of subcategories as filters, which narrows down the selection to a few candidates within the same subcategory. For final material selection, GPT-4V is prompted with detailed captions from the library, directing it to allocate the most suitable material. This is accomplished in conjunction with a meticulously engineered segmentation process in UV space, ultimately facilitating MLLM-based material matching. It is noteworthy that prompting GPT-4V to navigate through a three-level tree structure has been proven to enhance both the efficiency and accuracy of the matching process, as opposed to directly selecting from thousands of materials directly.
@@
B.5 Evaluation Details of Quantitative Results
Evaluation details. For the purpose of evaluating objects pre- and post-refinement, we render four view images of each object and concatenate them vertically to facilitate a comprehensive assessment. We craft a specific prompt to guide GPT-4V in conducting an impartial comparison of the texture quality between two objects. As demonstrated in Figure 14. GPT-4V’s advanced capabilities enable it to differentiate between the two objects, providing comparison results that closely align with assessments made by human experts. In the scenario evaluated, GPT-4V successfully identifies all enhancements (metal badges, human skin, and fabric textures).
Appendix C Additional Related Works
@@
Our problem setting poses a more challenging task, as illustrated in Figure 15, which involves identifying and recovering the original material properties from images of objects that only have a albedo map, in addition to generating different material maps for the entire object. Humans have the capacity to intuitively infer the underlying material properties of 3D objects from single images containing only shape and basic colors. This capability is attributed to our robust material recognition abilities and comprehensive prior knowledge of object categories, colors, and material attributes. Building on this concept, our approach harnesses the potent image recognition capabilities and prior knowledge inherent in large-scale multimodal language models to efficiently execute region-level material extraction and identification, which is further enhanced by subsequent 2D-3D alignment and albedo-referenced techniques to generate and apply material texture maps for physically realistic rendering of 3D objects. Furthermore,  [38] extracts the material takes 3 to 4 minutes, and process an image containing three materials takes more than 10 minutes. In contrast, our method takes only about 1 to 2 minutes to match and generate all the materials.
Figure 15: Comparison between previous method and Make-it-Real. We demonstrate the distinctions between Material Palette[38] and our method in terms of material identification and extraction. Our overall pipeline presents a more challenging task, where the input is a rendered image with only albedo information, and the output consists of textures for the entire object.
Appendix D Additional Experiments
D.1 Effects of Different Texture Maps